# HistoEncoder: a digital pathology foundation model for prostate cancer

### v1

@@ 2 Materials and methods
Non-overlapping $256\times 256$ tiles with ¿50% tissue; Part of HelsinkiProstate dataset
@@ 2.2.2 Combining histomics with other data modalities
In contrast, if the data modalities to be integrated contain information for larger tissue regions or whole or multiple slide images, the slide images must be cut into smaller tile images that are compatible with HistoEncoder. A common example is patient-level clinical data, where data from each patient pertains to all tile images from the same patient. In this case, we will first automatically annotate and cluster all tile images following the workflow in Sec. 2.2.1. Each cluster contains tile images with similar histological patterns. Second, we will calculate, for each patient, the proportion of tile images assigned to each cluster (”cluster fractions”). Cluster fractions represent a patient-specific summary of histological patterns observed in the slide images.
@@ 2.3 Datasets
All datasets leveraged in this study are summarized in Tab. 1. All slide images are processed with HistoPrep [46].
@@ 2.3.1 Training dataset
Training data, designated as HelsinkiProstate, for the prostate tissue encoders, consists of patients who have undergone prostate biopsies or radical prostatectomy in Helsinki University Hospital between 2013 and 2021. In total, there are 1,307 patients with 5,642 and 5,584 tissue slides of needle biopsies, and tissue sections from radical prostatectomy (Tab. 1). To create the HelsinkiProstate dataset, each slide is cut into $640\times 640$ pixel tile images at magnifications 20x, 10x, and 5x with 20% overlap between neighbouring tile images. This produces several hundred million tile images, which are then preprocessed with HistoPrep [46] to remove tiles containing non-tissue areas such as pen markings or other artefacts. Once irrelevant tile images have been filtered out, all remaining tile images are run through a prostate cancer classifier model from [47]. To create a balanced training dataset, all 16 million tile images with a prediction score $\hat{y}>0.2$, and 32 million randomly sampled tile images with score $\hat{y}\leq 0.2$ were selected to comprise the HelsinkiProstate dataset.
@@ 2.3.2 Evaluation datasets
Helsinki30 and Helsinki60 datasets [47] contain whole slide images (normal size and whole mounts) from 30 and 60 patients who have undergone radical prostatectomy at the Helsinki University Hospital between the years 2014 and 2021. All slide images in both datasets have been annotated by pathologists, and classified as cancerous or benign. These datasets are also part of the training (HelsinkiProstate) dataset (Tab. 1).
@@ 2.4.1 Prostate cancer-specific death survival models
To create survival models using HistoEncoder features, we first cluster the HelsinkiTMA dataset and obtain histological cluster memberships for each patient as described in Sec. 2.2.2 with the prostate-m encoder and number of clusters set to 32.
@@ 3.2 Identifying histological patterns as HistoEncoder feature clusters
To quantify whether similar features are extracted for tile images with similar histological patterns, the prostate-m encoder is used to automatically label the Karolinska and Radboud datasets with the annotation workflow (Sec. 2.2.1). After assigning each tile image to a cluster, the proportion of different labels in each feature cluster can be assessed. In the Karolinska dataset, 45.6% of all tile images, 53.3% of cancerous epithelium, and 42.4% of benign epithelium and stroma are contained in clusters with a label purity above 90% (Fig. 4). In the Radboud dataset, a clear majority of images belong to high-purity clusters (68.8% all tiles, 74.7% cancerous epithelium, and 63.9% benign epithelium and stroma). In total, 2.1 million $384\times 384$ tile images in the PANDA dataset could be classified into cancerous and benign tissue with over 90% accuracy after visually inspecting only 256 tile image clusters.
@@ 3.3 Predicing prostate cancer-specific mortality
We next combine information extracted with HistoEncoder from histological slide images with clinical data to build multimodal survival models predicting prostate cancer-specific death. The prostate-m encoder is used to collect patient-level feature cluster frequencies ($n=32$ clusters; Sec. 2.2.2) for the tissue microarray spots in the HelsinkiTMA dataset. These cluster frequencies represent the histological pattern distribution for each patient and can be simply concatenated with clinical data. Next, six feature clusters are selected and Cox proportional hazards models are fitted (see Sec. 2.4.1). A visual assessment of the top six feature clusters suggests more atypia in the cellular morphologies in clusters associated with worse prognoses (Supplementary Fig. 9).Certain subtypes and growth patterns, such as cribriform gland architecture and mucinous pattern are recurrent in the clusters with the highest HRs. Extracellular matrix formation, individual cell clusters and lymphocyte composition seem to vary even between the top six clusters.
@@ 4.1 Extending HistoEncoder
Supplementary Material
Table 2: AUROC scores for a K-nearest neighbor prostate cancer classifier, fitted with features extracted from the PESO dataset.
Model Parameters Training method Training data Karolinska Radboud Helsinki30 Helsinki60
XCiT-S12 25.9 self-supervised HelsinkiProstate 0.819 0.903 0.939 0.896
XCiT-M24 83.9 self-supervised HelsinkiProstate 0.824 0.898 0.943 0.901
XCiT-S12 25.9 self-supervised ImageNet 0.655 0.779 0.819 0.795
XCiT-M24 83.9 self-supervised ImageNet 0.683 0.790 0.816 0.806
EfficientNet-B0 4.0 supervised ImageNet 0.668 0.774 0.550 0.552
EfficientNet-B1 6.5 supervised ImageNet 0.666 0.753 0.545 0.54
EfficientNet-B2 7.7 supervised ImageNet 0.671 0.767 0.558 0.563
EfficientNet-B3 10.7 supervised ImageNet 0.661 0.740 0.549 0.548
EfficientNet-B4 17.5 supervised ImageNet 0.652 0.747 0.547 0.553
ResNet-18 11.2 supervised ImageNet 0.682 0.762 0.547 0.553
ResNet-50 23.5 supervised ImageNet 0.684 0.746 0.525 0.521
ResNet-101 42.5 supervised ImageNet 0.618 0.676 0.531 0.531
XCiT-S12 25.9 supervised ImageNet 0.645 0.690 0.707 0.714
XCiT-M24 83.9 supervised ImageNet 0.647 0.720 0.752 0.739
Table 3: AUROC scores for a K-nearest neighbor prostate cancer classifier, fitted with features extracted from the Karolinska dataset.
Model Parameters Training method Training data PESO Radboud Helsinki30 Helsinki60
XCiT-S12 25.9 self-supervised HelsinkiProstate 0.936 0.916 0.949 0.908
XCiT-M24 83.9 self-supervised HelsinkiProstate 0.940 0.918 0.945 0.902
XCiT-S12 25.9 self-supervised ImageNet 0.857 0.811 0.821 0.808
XCiT-M24 83.9 self-supervised ImageNet 0.837 0.790 0.796 0.790
EfficientNet-B0 4.0 supervised ImageNet 0.685 0.778 0.526 0.544
EfficientNet-B1 6.5 supervised ImageNet 0.707 0.778 0.529 0.551
EfficientNet-B2 7.7 supervised ImageNet 0.692 0.770 0.523 0.544
EfficientNet-B3 10.7 supervised ImageNet 0.744 0.784 0.541 0.551
EfficientNet-B4 17.5 supervised ImageNet 0.704 0.749 0.541 0.547
ResNet-18 11.2 supervised ImageNet 0.734 0.770 0.531 0.553
ResNet-50 23.5 supervised ImageNet 0.710 0.718 0.529 0.530
ResNet-101 42.5 supervised ImageNet 0.643 0.706 0.511 0.522
XCiT-S12 25.9 supervised ImageNet 0.784 0.683 0.664 0.648
XCiT-M24 83.9 supervised ImageNet 0.736 0.738 0.664 0.631
Table 4: AUROC scores for a K-Nearest neighbor prostate cancer classifier, fitted with features extracted from the Radboud dataset.
Model Parameters Training method Training data PESO Karolinska Helsinki30 Helsinki60
XCiT-S12 25.9 self-supervised HelsinkiProstate 0.934 0.854 0.959 0.919
XCiT-M24 83.9 self-supervised HelsinkiProstate 0.934 0.852 0.956 0.910
XCiT-S12 25.9 self-supervised ImageNet 0.875 0.757 0.831 0.818
XCiT-M24 83.9 self-supervised ImageNet 0.865 0.756 0.819 0.804
EfficientNet-B0 4.0 supervised ImageNet 0.768 0.740 0.545 0.545
EfficientNet-B1 6.5 supervised ImageNet 0.764 0.736 0.538 0.537
EfficientNet-B2 7.7 supervised ImageNet 0.745 0.721 0.537 0.537
EfficientNet-B3 10.7 supervised ImageNet 0.773 0.718 0.559 0.550
EfficientNet-B4 17.5 supervised ImageNet 0.676 0.653 0.499 0.501
ResNet-18 11.2 supervised ImageNet 0.786 0.737 0.567 0.569
ResNet-50 23.5 supervised ImageNet 0.780 0.711 0.551 0.543
ResNet-101 42.5 supervised ImageNet 0.690 0.683 0.517 0.526
XCiT-S12 25.9 supervised ImageNet 0.805 0.714 0.763 0.746
XCiT-M24 83.9 supervised ImageNet 0.797 0.723 0.776 0.732
Table 5: Cox proportional hazards model coefficients and p-values with Gleason grade and HistoEncoder feature cluster frequencies.
Variable coefficient exp(coefficient) SE(coefficient) CI, lower 95% CI, upper 95% $p$
GG1 (ref.)
GG2 2.01 7.45 0.83 0.39 3.63 0.02
GG3 1.19 3.29 0.88 -0.54 2.92 0.18
GG4 2.59 13.31 0.84 0.94 4.24 ¡0.005
GG5 1.91 6.74 0.95 0.04 3.78 0.05
Cluster 25 3.16 23.47 0.70 1.79 4.52 ¡0.005
Cluster 9 3.06 21.40 0.97 1.16 4.97 ¡0.005
Cluster 6 2.64 13.98 1.39 -0.09 5.36 0.06
Cluster 13 -1.78 0.17 2.60 -6.88 3.32 0.49
Cluster 21 2.41 11.09 1.07 0.32 4.49 0.02
Cluster 16 4.89 133.20 1.61 1.74 8.05 ¡0.005
Concordance=0.81; L1-ratio=0.5; penalizer=0.001; observations=398; events=52
Table 6: Cox proportional hazards model coefficients and $p$-values with CAPRA-S nomogram and HistoEncoder feature cluster frequencies.
Variable coefficient exp(coefficient) SE(coefficient) CI, lower 95% CI, upper 95% $p$
CAPRA-S 0.45 1.58 0.11 0.24 0.67 ¡0.005
Cluster 25 0.60 1.82 0.13 0.34 0.85 ¡0.005
Cluster 9 0.43 1.54 0.11 0.21 0.66 ¡0.005
Cluster 6 0.34 1.40 0.12 0.10 0.58 0.01
Cluster 13 0.07 1.07 0.27 -0.46 0.59 0.80
Cluster 21 0.18 1.20 0.14 -0.09 0.45 0.19
Cluster 16 0.33 1.39 0.34 -0.34 1.00 0.33
Concordance=0.89; L1-ratio=0.5; penalizer=0.001; observations=285; events=26
Table 7: Cox proportional hazards model coefficients and $p$-values with MSKCC-S nomogram and feature cluster frequencies.
Variable coefficient standard deviation CI, lower 95% CI, upper 95% $p$
MSKCC-S -0.54 0.58 0.14 -0.82 -0.26 ¡0.005
Cluster 25 0.57 1.76 0.12 0.32 0.81 ¡0.005
Cluster 9 0.42 1.53 0.11 0.20 0.65 ¡0.005
Cluster 6 0.40 1.48 0.12 0.16 0.63 ¡0.005
Cluster 13 0.01 1.01 0.26 -0.50 0.52 0.98
Cluster 21 0.17 1.19 0.13 -0.09 0.43 0.20
Cluster 16 0.30 1.35 0.36 -0.40 1.00 0.40
Concordance=0.86; L1-ratio=0.5; penalizer=0.001; observations=285; events=26
(a) Limiting the number of fine-tuned encoder blocks.
(b) Limiting the amount of training data.
Figure 6: AUROC scores for the prostate cancer classifiers fine-tuned from prostate-s and natural-s encoder models using the Karolinska and Radboud datasets. Prostate-s based models achieve higher AUROC scores than natural-s based models with less fine-tuned parameters (a) and training data (b). A KNN-classifier achieves similar or significantly better performance than fully fine-tuned natural-s based models. When limiting the number of tile images (b), prostate-s based models significantly outperform natural-s based models even trained with 1000 times less data.
Figure 7: Kaplan-Meier survival curve estimations for Gleason grades and clinical nomograms in HelsinkiTMA cohort used for the prostate cancer death survival models.
Figure 8: Cluster purities for the Radboud dataset with a varying number of clusters.
Figure 9: Randomly sampled tile images from the clusters used in the prostate cancer death survival models.

### v2

@@ Abstract
11footnotetext: These authors contributed equally.
    $\dagger$ Correspondence:
      esa.pitkanen@helsinki.fi, tuomas.mirtti@helsinki.fi
      Code: https://github.com/jopo666/HistoEncoder
@@ 2 Materials and methods
Non-overlapping $256\times 256$ tiles with >50% tissue; Part of HelsinkiProstate dataset
@@ 2.2.2 Combining histomics with other data modalities
In contrast, if the data modalities to be integrated contain information for larger tissue regions or whole or multiple slide images, the slide images must be cut into smaller tile images that are compatible with HistoEncoder. A common example is patient-level clinical data, where data from each patient pertains to all tile images from the same patient. In this case, we will first automatically annotate and cluster all tile images following the workflow in section 2.2.1. Each cluster contains tile images with similar histological patterns. Second, we will calculate, for each patient, the proportion of tile images assigned to each cluster ("cluster fractions"). Cluster fractions represent a patient-specific summary of histological patterns observed in the slide images.
@@ 2.3 Datasets
All datasets leveraged in this study are summarized in table 1. All slide images are processed with HistoPrep [44].
@@ 2.3.1 Training dataset
Training data, designated as HelsinkiProstate, for the prostate tissue encoders, consists of patients who have undergone prostate biopsies or radical prostatectomy in Helsinki University Hospital between 2013 and 2021. In total, there are 1,307 patients with 5,642 and 5,584 tissue slides of needle biopsies, and tissue sections from radical prostatectomy (table 1). To create the HelsinkiProstate dataset, each slide is cut into $640\times 640$ pixel tile images at magnifications 20x, 10x, and 5x with 20% overlap between neighbouring tile images. This produces several hundred million tile images, which are then preprocessed with HistoPrep [44] to remove tiles containing non-tissue areas such as pen markings or other artefacts. Once irrelevant tile images have been filtered out, all remaining tile images are run through a prostate cancer classifier model from [13]. To create a balanced training dataset, all 16 million tile images with a prediction score $\hat{y}>0.2$, and 32 million randomly sampled tile images with score $\hat{y}\leq 0.2$ were selected to comprise the HelsinkiProstate dataset.
@@ 2.3.2 Evaluation datasets
Helsinki30 and Helsinki60 datasets [13] contain whole slide images (normal size and whole mounts) from 30 and 60 patients who have undergone radical prostatectomy at the Helsinki University Hospital between the years 2014 and 2021. All slide images in both datasets have been annotated by pathologists, and classified as cancerous or benign. These datasets are also part of the training (HelsinkiProstate) dataset (table 1).
@@ 2.4.1 Prostate cancer-specific death survival models
To create survival models using HistoEncoder features, we first cluster the HelsinkiTMA dataset and obtain histological cluster memberships for each patient as described in section 2.2.2 with the prostate-m encoder and number of clusters set to 32.
@@ 3.2 Identifying histological patterns as HistoEncoder feature clusters
To quantify whether similar features are extracted for tile images with similar histological patterns, the prostate-m encoder is used to automatically label the Karolinska and Radboud datasets with the annotation workflow (section 2.2.1). After assigning each tile image to a cluster, the proportion of different labels in each feature cluster can be assessed. In the Karolinska dataset, 45.6% of all tile images, 53.3% of cancerous epithelium, and 42.4% of benign epithelium and stroma are contained in clusters with a label purity above 90% (fig. 4). In the Radboud dataset, a clear majority of images belong to high-purity clusters (68.8% all tiles, 74.7% cancerous epithelium, and 63.9% benign epithelium and stroma). In total, 2.1 million $384\times 384$ tile images in the PANDA dataset could be classified into cancerous and benign tissue with over 90% accuracy after visually inspecting only 256 tile image clusters.
@@ 3.3 Predicing prostate cancer-specific mortality
We next combine information extracted with HistoEncoder from histological slide images with clinical data to build multimodal survival models predicting prostate cancer-specific death. The prostate-m encoder is used to collect patient-level feature cluster frequencies ($n=32$ clusters; section 2.2.2) for the tissue microarray spots in the HelsinkiTMA dataset. These cluster frequencies represent the histological pattern distribution for each patient and can be simply concatenated with clinical data. Next, six feature clusters are selected and Cox proportional hazards models are fitted (see section 2.4.1). A visual assessment of the top six feature clusters suggests more atypia in the cellular morphologies in clusters associated with worse prognoses (Supplementary Fig. 9).Certain subtypes and growth patterns, such as cribriform gland architecture and mucinous pattern are recurrent in the clusters with the highest HRs. Extracellular matrix formation, individual cell clusters and lymphocyte composition seem to vary even between the top six clusters.
@@ 4.1 Extending HistoEncoder
Acknowledgements
This work was supported by Cancer Foundation Finland [304667, 191118], Jane and Aatos Erkko Foundation [290520], Research Council of Finland [322675] and Hospital District of Helsinki and Uusimaa [TYH2018214, TYH2018222, TYH2019235, TYH2019249]. The authors also wish to acknowledge FIMM Digital Microscopy and Molecular Pathology Unit supported by HiLIFE and Biocenter Finland for imaging services, and CSC – IT Center for Science, Finland for generous computational resources on the LUMI supercomputer (LUMI Extreme Scale project).
Ethics statement
Ethical approvals for the use of human tissue material and clinicopathologic data were obtained from the Institutional Ethics Committee of Hospital District of Helsinki and Uusimaa (§70/16.5.2018; HUS/419/2018) and by the National Supervisory Authority for Welfare and Health (VALVIRA, D:no V/38176/2018). According to the national and European Union legislation on noninterventional medical research, the study was conducted without informed individual patient consent by permission of the Hospital District of Helsinki and Uusimaa (§105/21.12.2018; HUS/419/2018). The experiments conformed to the principles set out in the WMA Declaration of Helsinki and the Department of Health and Human Services Belmont Report.
Data availability
The original human subject and sample-related data is available upon request, provided the institutional ethical approval and research permit allow data sharing according to the particular request. Restrictions may apply to the availability of the internal Helsinki University Hospital datasets, which cannot be made publicly available due to general data protection regulations, national legislation and institutional guidelines. Publicly available datasets can be accessed through their respective publications. For data inquiries, please contact tuomas.mirtti@helsinki.fi.
Declaration of competing interests
The authors have no interests to declare.